Alex and Bess have been in partnership for many years. The partners, who share profits and losses on a 60:40 basis, respectively, wish to retire and have agreed to liquidate the business. Liquidation expenses are estimated to be $9,000. At the date the partnership ceases operations, the balance sheet is as follows:








Cash$58,000Liabilities$44,000
Noncash assets
170,000Alex, capital
102,000



Bess, capital
82,000
Total assets$228,000Total liabilities and capital$228,000


Part A: 

Prepare journal entries for the following transactions:

  1. Distributed safe cash payments to the partners.

  2. Paid $26,400 of the partnership’s liabilities.

  3. Sold noncash assets for $184,000.

  4. Distributed safe cash payments to the partners.

  5. Paid remaining partnership liabilities of $17,600.

  6. Paid $7,200 in liquidation expenses; no further expenses will be incurred.

  7. Distributed remaining cash held by the business to the partners.


Part B: 

Prepare a final statement of partnership liquidation.







ALEX AND BESS PARTNERSHIP
Statement of Partnership Liquidation

CashNon-cash AssetsLiabilitiesAlex, CapitalBess, Capital
Beginning balances




Distribution to partners




Paid liabilities




Sold noncash assets




Updated balances




Distribution to partners




Updated balances




Paid liabilities




Paid liquidation expenses




Updated balances




Distribution to partners




Closing balances
Answer #2
Answer to Part A
Journal entries
Debit Credit
1) Cash          184,000
To Non cash assets          170,000
To gain on sales of assets            14,000
(Being assets sold)
2) Gain on realization            14,000
To Alex capital (60%)              8,400
To Bess capital (40%)              5,600
(Being gain on realization allocated among partners)
3) Liabilities            26,400
To cash            26,400
(Being liabilities paid off)
4) Liabilities            17,600
To cash            17,600
(Being liabilities paid off)
5) Liquidation expense               7,200
To cash              7,200
(Being Liquidation expense Paid)
6) Alex capital               4,320
Bess capital               2,880
To liquidation expense              7,200
(Being Liquidation expense allocated to partners in their remaining capital ratio)
7) Alex capital          106,080
Bess capital            84,720
To cash          190,800
(Being distribution of remaining cash among partners)
Answer to Part B
Statement of Partnership liquidation
Cash + Non cash assets = Liabilities + Capital
Alex (60%) Bess (40%)
Balances before realization 58000 170000 44000 102000 82000
Sale of assets and division of Profit (Refer W.N -1) 184000 -170000 8400 5600
Balances after realization 242000 0 44000 110400 87600
Payment of Liabilities -44000 -44000
Balances after payment of liabilities 198000 0 0 110400 87600
Payment of Liquidation expense -7200 -4320 -2880
Balances after payment of Liquidation expense 190800 0 0 106080 84720
Cash distributed to Partners -190800 -106080 -84720
Final balances 0 0 0 0 0 0
W.N - 1 ) Profit on realisation of Non cash assets
Book value of Assets $170,000
Realisation from sale of assets $184,000
Profit on sale of assets $14,000
Answer #1

Alex and Bess Partnership

Statement of Partnership Liquidation

Cash

Non-cash assets

Liabilities

Alex capital (60%)

Bess capital (40%)

Beginning balances

58000

170000

(44000)

(102000)

(82000)

Distribution to partners

(5000)

5000

Paid liabilities

(26400)

26400

Sold noncash assets

184000

(170000)

(8400)

(5600)

Updated balances

210600

0

(17600)

(110400)

(82600)

Distribution to partners

(184000)

105252

78748

Updated balances

26600

0

(17600)

(5148)

(3852)

Paid liabilities

(17600)

17600

Paid liquidation expenses

(7200)

4119

3081

Updated balances

1800

0

0

(1029)

(771)

Distribution to partners

(1800)

1029

771

Closing balances

0

0

0

0

0

Estimated cash after setting of creditors and liquidation Expenses = 58000-44000-9000 = 5000

Alex

Bess

Capital

102000

82000

Assumed losses (nets assets)

(102000)

(68000)

Sub total

0

14000

Safe payments

5000

Alex = 184000*110400/(110400+82600) = 105252

Bess =184000*82600/(110400+82600) = 74748

Alex = 7200*110400/(110400+82600) = 105252

Bess =7200*82600/(110400+82600) = 74748

Alex = 1800*110400/(110400+82600) = 1029

Bess =1800*82600/(110400+82600) = 771